# Strong Linearizability without Compare&Swap: The Case of Bags

### v2

@@ 7.1 Strong-Linearizability
    We prove in Lemma 14 below that at most one successful t&s() can be performed on TS[a] after insins writes to Items[a] and before it writes to InsertDone. Hence, an Insert operation is coupled with at most one Take operation.
@@
Proof 10
Before the producer writes an element $x$ to Items[m] on LABEL:b-bounded:insert:write_item, it adds m to alloc on LABEL:b-bounded:insert:add_to_alloc. Immediately after the producer writes $\bot$ to Items[m] on LABEL:b-bounded:insert:clear_item, it removes m from alloc on LABEL:b-bounded:insert:remove_from_alloc.
@@
Proof 11
The producer adds the location m to used on LABEL:b-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:b-bounded:insert:clear_item. When the producer writes an element $x$ to Items[m] on LABEL:b-bounded:insert:write_item, m is not in used. This is because used is a subset of hazardous, from LABEL:b-bounded:insert:narrow_used_down, and m s not in hazardous, from LABEL:b-bounded:insert:choose_index.
@@
Note that TS[$a$] = 0 immediately before $tk$ performed its successful t&s() on TS[$a$]. Let $s_{0}$ be the step when the producer last wrote to Items[$a$] (on LABEL:b-bounded:insert:write_item) prior to $s_{1}$. So, Items[$a$] = $x$ between $s_{0}$ and $s_{1}$. Assume, for the sake of contradiction, that TS[$a$] = 1 at some configuration between $s_{0}$ and $s_{2}$. Then between this configuration and $s_{2}$, the producer must have reset TS[$a$] on LABEL:b-bounded:insert:reset_TS. From LABEL:b-bounded:insert:reset_loop, $a$ was in used when the reset was performed, so, by Lemma 11, Items[$a$] = $\bot$ when the reset was performed. Thus, the producer must have written $\bot$ to Items[$a$] on LABEL:b-bounded:insert:clear_item before it performed the reset. Since Items[$a$] =x=x between $s_{0}$ and $s_{1}$, the producer must have written $\bot$ to Items[$a$], collected Hazards on LABEL:b-bounded:insert:collect_hazards, and performed the reset between $s_{1}$ and $s_{2}$. In particular, the producer reads $a$ from Hazards[i]. But, by LABEL:b-bounded:insert:reset_loop, the producer does not reset TS[$a$] on LABEL:b-bounded:insert:reset_TS. This is a contradiction. Thus TS[$a$] = 0 between $s_{0}$ and $s_{2}$.
By the test on LABEL:b-bounded:insert:check_TS, the producer does not write $\bot$ to Items[$a$] on LABEL:b-bounded:insert:clear_item and $a$ is not removed from alloc on LABEL:b-bounded:insert:remove_from_alloc between $s_{0}$ and $s_{2}$. Thus, the producer does not choose the index $a$ again on LABEL:b-bounded:insert:choose_index and does not write to Items[$a$] on LABEL:b-bounded:insert:write_item between these two steps. Hence, Items[$a$] = $x$ between when the producer last wrote to Items[$a$] at step $s_{0}$ and when $tk$ performed its successful t&s() at step $s_{2}$.
@@
By Lemma 11, Items[a] = $\bot$ while $a\in$ used. Immediately after step $s$, Items[$a$] $\neq\bot$, so $a\not\in$ used. Prior to being removed from used on LABEL:b-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:b-bounded:insert:reset_TS. Since $a\in$ used when this step occurred, Items[$a$] = $\bot$. Between this step and $s$, there is no write to Items[$a$] on LABEL:b-bounded:insert:write_item, by definition of $s^{\prime}$, so Items[$a$] = $\bot$. When a successful t&s() on TS[$a$] is performed, Lemma 12 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s() on TS[$a$] was performed between these two steps. Hence, TS[$a$] = 0 when $s$ was performed.
@@
Proof 14
By Lemma 13, when insins writes to Items[a] on LABEL:b-bounded:insert:write_item, TS[a] = 0. Assume some Take operation performs a successful t&s() on TS[a] after insins performs this write and before it performs its next step, which is a write to InsertDone on LABEL:b-bounded:insert:write_InsertDone. For another Take operation to perform a successful t&s() on TS[a], TS[a] must be first reset to 0. It is reset only by the producer (on LABEL:b-bounded:insert:reset_TS). Therefore, no other Take operation can perform a successful t&s() on TS[a] while insins is poised to write to InsertDone.
@@
Proof 15
Let $ts$ be a successful TS[$a$].t&s() performed by a Take operation on LABEL:b-bounded:take:t&s. By Lemma 12, when $ts$ is performed, Items[$a$] $\neq\bot$ and, hence, by Lemma 11, $a\notin$ used. When a reset of TS[$a$] is performed on LABEL:b-bounded:insert:reset_TS, the condition on LABEL:b-bounded:insert:reset_loop ensures that $a\in$ used. So $a$ is added to used on LABEL:b-bounded:insert:add_to_used between $ts$ and the next reset of TS[$a$]. Before $a$ is added to used, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS.
Only the producer performs resets, but does not perform any resets between performing LABEL:b-bounded:insert:check_TS and next performing LABEL:b-bounded:insert:add_to_used. Since TS[a] =0=0 when $ts$ is performed, $ts$ does not occur between these two steps by the producer. Hence, the producer reads 1 from TS[a] after $ts$ and before it next resets TS[$a$].
@@
Proof 16
Suppose that $tk_{1}$ is a Take operation that performs a successful t&s(), $ts_{1}$, on TS[$a$] before the Take operation $tk_{2}$ performs a successful t&s(), $ts_{2}$, on TS[$a$]. A reset of TS[$a$] occurs between $ts_{1}$ and $ts_{2}$. By Lemma 15, the producer reads 1 from TS[a] (on LABEL:b-bounded:insert:check_TS) between $ts_{1}$ and performing this reset. Note that $tk_{1}$ is linearized at or before this read. However, $tk_{2}$ is linearized after $ts_{2}$. Thus, they are not linearized at the same step.
@@
Proof 17
Let $tk$ be the Take operation that performs the first successful TS[a].t&s(), $ts$, following the write, $w_{1}$, of $x_{1}$ to Items[$a$]. Let $r$ be the last read of Items[a] on LABEL:b-bounded:take:read_item performed by $tk$ prior to $ts$ and let $x$ be the element that it read during $r$. Let $w$ be the last write to Items[a] on LABEL:b-bounded:insert:write_item by the producer before $r$, which implies this is a write of $x$. Between $w$ and the producer’s next write to Items[a], the producer executes LABEL:b-bounded:insert:choose_index, and at that point $a\notin$ alloc, which implies by Lemma 10 that Items[a] = $\bot$. But by Lemma 12, Items[a] = $x$ between $w$ and $ts$. Therefore, the producer’s next write to Items[a] after $w$ must happen after $ts$. In other words, $w$ was the last write to Items[a] before $ts$, namely, $w=w_{1}$. Hence, $x=x_{1}$. So $tk$ would return $x_{1}$.
@@
Let $S$ be the ordering of a prefix, $\sigma$, of the execution, let $e$ be the next step of the execution, and $S^{\prime}$ be the ordering of $\sigma\cdot e$. Suppose $S$ is a linearization. Let $B$ be the set of items in the bag at the end of $S$. Further suppose that for every Take operation $tk$ that appears in $S$, returns an item $x$, and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$, an Insert($x$) operation that wrote in $\sigma$ to Items[a] appears in $S$ before $tk$ and no Take operation returning $x$ appear in $S$ between this Insert operation and $tk$. Due to our no repetitions assumption, this informally means that every Take operation that performs in $\sigma$ a successful t&s() on some location in TS, takes the item that was inserted to the bag by an Insert operation that wrote to the same location in Items. Let $B$ be the set of items in the bag at the end of $S$. We will prove that $S^{\prime}$ is a linearization, and that for every Take operation $tk$ that appears in $S^{\prime}$ but not in $S$, returns an item $x$ and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$, an Insert($x$) operation that wrote in $\sigma$ to Items[a] appears in $S$, such that no Take operation returning $x$ appears in $S$ after it. The latter part is proven in Lemma 21.
@@
Proof 18
Let $ins_{2}^{\prime}$ be the first successful Insert operation to be executed after $ins_{1}$ among those allocated location $a$. Let $w_{1}$ be the write event to Items[a] by $ins_{1}$ on LABEL:b-bounded:insert:write_item, and $w_{2}$ be the write event to Items[a] by $ins_{2}^{\prime}$. Let $w_{InsertDone}$ be the write event to InsertDone by $ins_{1}$ on LABEL:b-bounded:insert:write_InsertDone, which is its next step after writing to Items[a]. From the code, it is evident that $a\in$ alloc when $w_{InsertDone}$ occurs. Before $ins_{2}^{\prime}$ writes to Items[a], it allocates location $a$ to m on LABEL:b-bounded:insert:choose_index, and at that point $a\notin$ alloc. Therefore, between $w_{InsertDone}$ and $w_{2}$, $m$ is removed from alloc. This is done by the producer on LABEL:b-bounded:insert:remove_from_alloc, after reading 1 from TS[a] on LABEL:b-bounded:insert:check_TS. Let $r_{TS}$ be this read event. Then $r_{TS}$ occurs between $w_{InsertDone}$ and $w_{2}$. By Lemma 13, TS[a] = 0 at $w_{1}$. Thus, between $w_{1}$ and $r_{TS}$, a Take operation performed a successful TS[a].t&s(). Let $tk$ be the Take operation to perform the first such t&s(), and let $ts$ be this t&s() event. $tk$ returns the item inserted by $ins_{1}$ due to Lemma 17.
@@
Proof 19
By Lemma 13, TS[a] = 0 at $w$. Hence, between $ts$ and $w$, a reset of TS[a] on LABEL:b-bounded:insert:reset_TS must take place. By Lemma 15, between $ts$ and this reset, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS. $tk$ is linearized at this read if it is not already linearized earlier. Therefore, $tk$ is linearized before $w$, and thus, before insins.
@@
Proof 20
Let $ts$ be $tk$’s successful TS[a].t&s(). Before another successful TS[a].t&s() occurs in $\sigma$ after $ts$, a reset of TS[a] must be executed. By Lemma 15, between $ts$ and this reset, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS. But no such read is executed in $\sigma$ after $ts$, because if it were, $tk$ would be linearized at this step at the latest and would appear in $S$. is executed in $\sigma$ after $ts$. We have that a successful TS[a].t&s() that follows $ts$ must occur after a reset, which in turn occurs after read of 1 from TS[a], which does not occur in $\sigma$ after $ts$. So such a successful t&s() does not occur in $\sigma$.
@@
Proof 21
Let $ts$ be $tk$’s successful TS[a].t&s(). Let $w$ be the write to Items[a] on LABEL:b-bounded:insert:write_item by insins. We first consider Take operations in $S$ that perform in $\sigma$ a successful t&s() on TS[a], if such operations exist. No successful t&s() is performed on TS[a] between $w$ and $ts$, as TS[a] = 0 in this interval by Lemma 12. No successful t&s() is performed on TS[a] in $\sigma$ after $ts$ by Lemma 20. Any Take operation performing a successful TS[a].t&s() before $w$, appears in $S$ before insins by Lemma 19.
Now, we consider Take operations in $S$ returning $x$, which perform in $\sigma$ a successful t&s() on TS[a’] for some index $a^{\prime}\neq a$. We will show that if any such operations exist, they must appear before insins in $S$. Assume for sake of contradiction that such an operation appears after insins in $S$. Let $tk^{\prime}$ be the first such operation to appear in $S$ after insins. Assume it performed a successful t&s() on TS[a’] for some location $a^{\prime}$. Due to the induction hypothesis, an Insert($x$) operation that wrote in $\sigma$ to Items[a’] appears in $S$ before $tk^{\prime}$, such that $tk^{\prime}$ is the first Take operation returning $x$ to appear after this Insert($x$). This Insert operation is not insins, because they write to different locations in Items. $tk^{\prime}$ is also the first Take operation returning $x$ to appear in $S$ after insins, due to the definition of $tk^{\prime}$ in addition to what we earlier proved—that Take operations that perform in $\sigma$ a successful t&s() on TS[a] do not appear in $S$ after insins. Hence, right before $tk^{\prime}$, $x$ appears at least twice in the bag. This is a contradiction to our no repetitions assumption.
@@
Proof 22
alloc is initialized to $\phi$. A location is added to alloc only on LABEL:b-bounded:insert:add_to_alloc, when $|alloc|<b$.
@@
$|alloc|=b$ when insins executes LABEL:b-bounded:insert:alloc<b after $C^{*}$: $|alloc|$ cannot be smaller than $b$ at that step because insins continues to the else if clause, and it cannot be bigger than $b$ due to Lemma 22. insins does not execute LABEL:b-bounded:insert:remove_from_alloc between $C^{*}$ and $C$, because remove a location from alloc would imply that $|alloc|>b$ before the removal, which is impossible by Lemma 22. So insins does not modify allocalloc between $C^{*}$ and $C$. Let $alloc^{*}$ be the value of allocalloc in this interval. We have $|alloc^{*}|=b$. This implies that insins obtains 0 from the read of TS[m] for all indices $m\in alloc$ in its executions of LABEL:b-bounded:insert:check_TS following $C^{*}$.
@@
Let $w$ be the write to Items[m] on LABEL:b-bounded:insert:write_item by $ins^{\prime}$. At $w$, $m\notin$ hazardous according to LABEL:b-bounded:insert:choose_index, while used $\subseteq$ hazardous by LABEL:b-bounded:insert:narrow_used_down. Therefore, $m$ $\notin$ used at $w$.After $w$, $m$ is added to used on LABEL:b-bounded:insert:add_to_used only after it is removed from alloc on LABEL:b-bounded:insert:remove_from_alloc. The latter may happen only after $C$: it cannot happen until $C^{*}$ by definition of $ins^{\prime}$, and it cannot happen between $C^{*}$ until $C$ since alloc remains unchanged in this interval. Thus, $m$ $\notin$ used continues to hold from $w$ until $C$. Therefore, no reset of TS[m] is executed on LABEL:b-bounded:insert:reset_TS between $w$ and $C$. Recall that $ins^{\prime}$ reads TS[m] between $C^{*}$ and $C$ and obtains 0. As TS[m] is not reset between $w$ and this read, a successful TS[m].t&s() cannot be executed within this interval, specifically not between $w$ and $C^{*}$. A Take operation that executes a successful TS[m].t&s() before $w$ is not linearized after $ins^{\prime}$ by Lemma 19, and a Take operation that executes a successful TS[m].t&s() after $C^{*}$ is not linearized at $S^{*}$. Therefore, no Take operation is linearized after $ins^{\prime}$ in $S^{*}$.
Next, suppose $e$ is an execution of a read of 1 from TS[a] by an Insert operation insins on LABEL:b-bounded:insert:check_TS. Further suppose a Take operation, $tk$, performed a successful t&s() on TS[a] on LABEL:b-bounded:take:t&s, in a step $ts$. Assume between $ts$ and $e$, none of the following occurred: another read of 1 from TS[a] by the producer, a write to TakeDone, or a write to Items[a’] for $a^{\prime}\neq a$ by an uncoupled Insert operation. There could not be more than one such Take operation by Lemma 16. Let the item $tk$ obtained from Items[a] in its last execution of LABEL:b-bounded:take:read_item be $x$. Then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $x\in B$.
Let $ins^{\prime}$ be the last Insert operation to write to Items[a] on LABEL:b-bounded:insert:write_item before $tk$ reads it for the last time (which it does before $ts$), and let $w$ be this write. Then $w$ is a write of $x$ and $ins^{\prime}$ in an Insert($x$) operation. $ins^{\prime}\neq ins$ since $ins^{\prime}$ does not execute LABEL:b-bounded:insert:check_TS after $w$. Therefore, $ins^{\prime}$ is completed in $\sigma$ and thus appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins^{\prime}$, the bag contains $x$. By Lemma 21, there is no Take operation that returns $x$ in $S$ after $ins^{\prime}$. This implies $x\in B$.

### v3

@@ 7.1 Strong-Linearizability
    We prove in Lemma 7.5 below that at most one successful t&s() can be performed on TS[a] after insins writes to Items[a] and before it writes to InsertDone. Hence, an Insert operation is coupled with at most one Take operation.
@@
Proof.
Before the producer writes an element $x$ to Items[m] on LABEL:b-bounded:insert:write_item, it adds m to alloc on LABEL:b-bounded:insert:add_to_alloc. Immediately after the producer writes $\bot$ to Items[m] on LABEL:b-bounded:insert:clear_item, it removes m from alloc on LABEL:b-bounded:insert:remove_from_alloc. ∎
@@
Proof.
The producer adds the location m to used on LABEL:b-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:b-bounded:insert:clear_item. When the producer writes an element $x$ to Items[m] on LABEL:b-bounded:insert:write_item, m is not in used. This is because used is a subset of hazardous, from LABEL:b-bounded:insert:narrow_used_down, and m s not in hazardous, from LABEL:b-bounded:insert:choose_index. ∎
@@
Note that TS[$a$] = 0 immediately before $tk$ performed its successful t&s() on TS[$a$]. Let $s_{0}$ be the step when the producer last wrote to Items[$a$] (on LABEL:b-bounded:insert:write_item) prior to $s_{1}$. So, Items[$a$] = $x$ between $s_{0}$ and $s_{1}$. Assume, for the sake of contradiction, that TS[$a$] = 1 at some configuration between $s_{0}$ and $s_{2}$. Then between this configuration and $s_{2}$, the producer must have reset TS[$a$] on LABEL:b-bounded:insert:reset_TS. From LABEL:b-bounded:insert:reset_loop, $a$ was in used when the reset was performed, so, by Lemma 7.2, Items[$a$] = $\bot$ when the reset was performed. Thus, the producer must have written $\bot$ to Items[$a$] on LABEL:b-bounded:insert:clear_item before it performed the reset. Since Items[$a$] =x=x between $s_{0}$ and $s_{1}$, the producer must have written $\bot$ to Items[$a$], collected Hazards on LABEL:b-bounded:insert:collect_hazards, and performed the reset between $s_{1}$ and $s_{2}$. In particular, the producer reads $a$ from Hazards[i]. But, by LABEL:b-bounded:insert:reset_loop, the producer does not reset TS[$a$] on LABEL:b-bounded:insert:reset_TS. This is a contradiction. Thus TS[$a$] = 0 between $s_{0}$ and $s_{2}$.
By the test on LABEL:b-bounded:insert:check_TS, the producer does not write $\bot$ to Items[$a$] on LABEL:b-bounded:insert:clear_item and $a$ is not removed from alloc on LABEL:b-bounded:insert:remove_from_alloc between $s_{0}$ and $s_{2}$. Thus, the producer does not choose the index $a$ again on LABEL:b-bounded:insert:choose_index and does not write to Items[$a$] on LABEL:b-bounded:insert:write_item between these two steps. Hence, Items[$a$] = $x$ between when the producer last wrote to Items[$a$] at step $s_{0}$ and when $tk$ performed its successful t&s() at step $s_{2}$. ∎
@@
By Lemma 7.2, Items[a] = $\bot$ while $a\in$ used. Immediately after step $s$, Items[$a$] $\neq\bot$, so $a\not\in$ used. Prior to being removed from used on LABEL:b-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:b-bounded:insert:reset_TS. Since $a\in$ used when this step occurred, Items[$a$] = $\bot$. Between this step and $s$, there is no write to Items[$a$] on LABEL:b-bounded:insert:write_item, by definition of $s^{\prime}$, so Items[$a$] = $\bot$. When a successful t&s() on TS[$a$] is performed, Lemma 7.3 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s() on TS[$a$] was performed between these two steps. Hence, TS[$a$] = 0 when $s$ was performed. ∎
@@
Proof.
By Lemma 7.4, when insins writes to Items[a] on LABEL:b-bounded:insert:write_item, TS[a] = 0. Assume some Take operation performs a successful t&s() on TS[a] after insins performs this write and before it performs its next step, which is a write to InsertDone on LABEL:b-bounded:insert:write_InsertDone. For another Take operation to perform a successful t&s() on TS[a], TS[a] must be first reset to 0. It is reset only by the producer (on LABEL:b-bounded:insert:reset_TS). Therefore, no other Take operation can perform a successful t&s() on TS[a] while insins is poised to write to InsertDone. ∎
@@
Proof.
Let $ts$ be a successful TS[$a$].t&s() performed by a Take operation on LABEL:b-bounded:take:t&s. By Lemma 7.3, when $ts$ is performed, Items[$a$] $\neq\bot$ and, hence, by Lemma 7.2, $a\notin$ used. When a reset of TS[$a$] is performed on LABEL:b-bounded:insert:reset_TS, the condition on LABEL:b-bounded:insert:reset_loop ensures that $a\in$ used. So $a$ is added to used on LABEL:b-bounded:insert:add_to_used between $ts$ and the next reset of TS[$a$]. Before $a$ is added to used, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS.
Only the producer performs resets, but does not perform any resets between performing LABEL:b-bounded:insert:check_TS and next performing LABEL:b-bounded:insert:add_to_used. Since TS[a] =0=0 when $ts$ is performed, $ts$ does not occur between these two steps by the producer. Hence, the producer reads 1 from TS[a] after $ts$ and before it next resets TS[$a$]. ∎
@@
Proof.
Suppose that $tk_{1}$ is a Take operation that performs a successful t&s(), $ts_{1}$, on TS[$a$] before the Take operation $tk_{2}$ performs a successful t&s(), $ts_{2}$, on TS[$a$]. A reset of TS[$a$] occurs between $ts_{1}$ and $ts_{2}$. By Lemma 7.6, the producer reads 1 from TS[a] (on LABEL:b-bounded:insert:check_TS) between $ts_{1}$ and performing this reset. Note that $tk_{1}$ is linearized at or before this read. However, $tk_{2}$ is linearized after $ts_{2}$. Thus, they are not linearized at the same step. ∎
@@
Proof.
Let $tk$ be the Take operation that performs the first successful TS[a].t&s(), $ts$, following the write, $w_{1}$, of $x_{1}$ to Items[$a$]. Let $r$ be the last read of Items[a] on LABEL:b-bounded:take:read_item performed by $tk$ prior to $ts$ and let $x$ be the element that it read during $r$. Let $w$ be the last write to Items[a] on LABEL:b-bounded:insert:write_item by the producer before $r$, which implies this is a write of $x$. Between $w$ and the producer’s next write to Items[a], the producer executes LABEL:b-bounded:insert:choose_index, and at that point $a\notin$ alloc, which implies by Lemma 7.1 that Items[a] = $\bot$. But by Lemma 7.3, Items[a] = $x$ between $w$ and $ts$. Therefore, the producer’s next write to Items[a] after $w$ must happen after $ts$. In other words, $w$ was the last write to Items[a] before $ts$, namely, $w=w_{1}$. Hence, $x=x_{1}$. So $tk$ would return $x_{1}$. ∎
@@
Let $S$ be the ordering of a prefix, $\sigma$, of the execution, let $e$ be the next step of the execution, and $S^{\prime}$ be the ordering of $\sigma\cdot e$. Suppose $S$ is a linearization. Let $B$ be the set of items in the bag at the end of $S$. Further suppose that for every Take operation $tk$ that appears in $S$, returns an item $x$, and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$, an Insert($x$) operation that wrote in $\sigma$ to Items[a] appears in $S$ before $tk$ and no Take operation returning $x$ appear in $S$ between this Insert operation and $tk$. Due to our no repetitions assumption, this informally means that every Take operation that performs in $\sigma$ a successful t&s() on some location in TS, takes the item that was inserted to the bag by an Insert operation that wrote to the same location in Items. Let $B$ be the set of items in the bag at the end of $S$. We will prove that $S^{\prime}$ is a linearization, and that for every Take operation $tk$ that appears in $S^{\prime}$ but not in $S$, returns an item $x$ and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$, an Insert($x$) operation that wrote in $\sigma$ to Items[a] appears in $S$, such that no Take operation returning $x$ appears in $S$ after it. The latter part is proven in Lemma 7.12.
@@
Proof.
Let $ins_{2}^{\prime}$ be the first successful Insert operation to be executed after $ins_{1}$ among those allocated location $a$. Let $w_{1}$ be the write event to Items[a] by $ins_{1}$ on LABEL:b-bounded:insert:write_item, and $w_{2}$ be the write event to Items[a] by $ins_{2}^{\prime}$. Let $w_{InsertDone}$ be the write event to InsertDone by $ins_{1}$ on LABEL:b-bounded:insert:write_InsertDone, which is its next step after writing to Items[a]. From the code, it is evident that $a\in$ alloc when $w_{InsertDone}$ occurs. Before $ins_{2}^{\prime}$ writes to Items[a], it allocates location $a$ to m on LABEL:b-bounded:insert:choose_index, and at that point $a\notin$ alloc. Therefore, between $w_{InsertDone}$ and $w_{2}$, $m$ is removed from alloc. This is done by the producer on LABEL:b-bounded:insert:remove_from_alloc, after reading 1 from TS[a] on LABEL:b-bounded:insert:check_TS. Let $r_{TS}$ be this read event. Then $r_{TS}$ occurs between $w_{InsertDone}$ and $w_{2}$. By Lemma 7.4, TS[a] = 0 at $w_{1}$. Thus, between $w_{1}$ and $r_{TS}$, a Take operation performed a successful TS[a].t&s(). Let $tk$ be the Take operation to perform the first such t&s(), and let $ts$ be this t&s() event. $tk$ returns the item inserted by $ins_{1}$ due to Lemma 7.8.
@@
Proof.
By Lemma 7.4, TS[a] = 0 at $w$. Hence, between $ts$ and $w$, a reset of TS[a] on LABEL:b-bounded:insert:reset_TS must take place. By Lemma 7.6, between $ts$ and this reset, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS. $tk$ is linearized at this read if it is not already linearized earlier. Therefore, $tk$ is linearized before $w$, and thus, before insins. ∎
@@
Proof.
Let $ts$ be $tk$’s successful TS[a].t&s(). Before another successful TS[a].t&s() occurs in $\sigma$ after $ts$, a reset of TS[a] must be executed. By Lemma 7.6, between $ts$ and this reset, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS. But no such read is executed in $\sigma$ after $ts$, because if it were, $tk$ would be linearized at this step at the latest and would appear in $S$. is executed in $\sigma$ after $ts$. We have that a successful TS[a].t&s() that follows $ts$ must occur after a reset, which in turn occurs after read of 1 from TS[a], which does not occur in $\sigma$ after $ts$. So such a successful t&s() does not occur in $\sigma$. ∎
@@
Proof.
Let $ts$ be $tk$’s successful TS[a].t&s(). Let $w$ be the write to Items[a] on LABEL:b-bounded:insert:write_item by insins. We first consider Take operations in $S$ that perform in $\sigma$ a successful t&s() on TS[a], if such operations exist. No successful t&s() is performed on TS[a] between $w$ and $ts$, as TS[a] = 0 in this interval by Lemma 7.3. No successful t&s() is performed on TS[a] in $\sigma$ after $ts$ by Lemma 7.11. Any Take operation performing a successful TS[a].t&s() before $w$, appears in $S$ before insins by Lemma 7.10.
Now, we consider Take operations in $S$ returning $x$, which perform in $\sigma$ a successful t&s() on TS[a’] for some index $a^{\prime}\neq a$. We will show that if any such operations exist, they must appear before insins in $S$. Assume for sake of contradiction that such an operation appears after insins in $S$. Let $tk^{\prime}$ be the first such operation to appear in $S$ after insins. Assume it performed a successful t&s() on TS[a’] for some location $a^{\prime}$. Due to the induction hypothesis, an Insert($x$) operation that wrote in $\sigma$ to Items[a’] appears in $S$ before $tk^{\prime}$, such that $tk^{\prime}$ is the first Take operation returning $x$ to appear after this Insert($x$). This Insert operation is not insins, because they write to different locations in Items. $tk^{\prime}$ is also the first Take operation returning $x$ to appear in $S$ after insins, due to the definition of $tk^{\prime}$ in addition to what we earlier proved—that Take operations that perform in $\sigma$ a successful t&s() on TS[a] do not appear in $S$ after insins. Hence, right before $tk^{\prime}$, $x$ appears at least twice in the bag. This is a contradiction to our no repetitions assumption. ∎
@@
Proof.
alloc is initialized to $\phi$. A location is added to alloc only on LABEL:b-bounded:insert:add_to_alloc, when $|alloc|<b$. ∎
@@
$|alloc|=b$ when insins executes LABEL:b-bounded:insert:alloc<b after $C^{*}$: $|alloc|$ cannot be smaller than $b$ at that step because insins continues to the else if clause, and it cannot be bigger than $b$ due to Lemma 7.13. insins does not execute LABEL:b-bounded:insert:remove_from_alloc between $C^{*}$ and $C$, because remove a location from alloc would imply that $|alloc|>b$ before the removal, which is impossible by Lemma 7.13. So insins does not modify allocalloc between $C^{*}$ and $C$. Let $alloc^{*}$ be the value of allocalloc in this interval. We have $|alloc^{*}|=b$. This implies that insins obtains 0 from the read of TS[m] for all indices $m\in alloc$ in its executions of LABEL:b-bounded:insert:check_TS following $C^{*}$.
@@
Let $w$ be the write to Items[m] on LABEL:b-bounded:insert:write_item by $ins^{\prime}$. At $w$, $m\notin$ hazardous according to LABEL:b-bounded:insert:choose_index, while used $\subseteq$ hazardous by LABEL:b-bounded:insert:narrow_used_down. Therefore, $m$ $\notin$ used at $w$.After $w$, $m$ is added to used on LABEL:b-bounded:insert:add_to_used only after it is removed from alloc on LABEL:b-bounded:insert:remove_from_alloc. The latter may happen only after $C$: it cannot happen until $C^{*}$ by definition of $ins^{\prime}$, and it cannot happen between $C^{*}$ until $C$ since alloc remains unchanged in this interval. Thus, $m$ $\notin$ used continues to hold from $w$ until $C$. Therefore, no reset of TS[m] is executed on LABEL:b-bounded:insert:reset_TS between $w$ and $C$. Recall that $ins^{\prime}$ reads TS[m] between $C^{*}$ and $C$ and obtains 0. As TS[m] is not reset between $w$ and this read, a successful TS[m].t&s() cannot be executed within this interval, specifically not between $w$ and $C^{*}$. A Take operation that executes a successful TS[m].t&s() before $w$ is not linearized after $ins^{\prime}$ by Lemma 7.10, and a Take operation that executes a successful TS[m].t&s() after $C^{*}$ is not linearized at $S^{*}$. Therefore, no Take operation is linearized after $ins^{\prime}$ in $S^{*}$.
Next, suppose $e$ is an execution of a read of 1 from TS[a] by an Insert operation insins on LABEL:b-bounded:insert:check_TS. Further suppose a Take operation, $tk$, performed a successful t&s() on TS[a] on LABEL:b-bounded:take:t&s, in a step $ts$. Assume between $ts$ and $e$, none of the following occurred: another read of 1 from TS[a] by the producer, a write to TakeDone, or a write to Items[a’] for $a^{\prime}\neq a$ by an uncoupled Insert operation. There could not be more than one such Take operation by Lemma 7.7. Let the item $tk$ obtained from Items[a] in its last execution of LABEL:b-bounded:take:read_item be $x$. Then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $x\in B$.
Let $ins^{\prime}$ be the last Insert operation to write to Items[a] on LABEL:b-bounded:insert:write_item before $tk$ reads it for the last time (which it does before $ts$), and let $w$ be this write. Then $w$ is a write of $x$ and $ins^{\prime}$ in an Insert($x$) operation. $ins^{\prime}\neq ins$ since $ins^{\prime}$ does not execute LABEL:b-bounded:insert:check_TS after $w$. Therefore, $ins^{\prime}$ is completed in $\sigma$ and thus appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins^{\prime}$, the bag contains $x$. By Lemma 7.12, there is no Take operation that returns $x$ in $S$ after $ins^{\prime}$. This implies $x\in B$.